### v2

@@ 3 Preliminaries
Semi-supervised Graph Node Classification. Graph node classification is a fundamental task in graph representation learning, with the purpose of classifying graph nodes into predefined categories [53], and can be used as a proxy task for measuring the discrimination ability of graph representation models. Existing methods mainly focus on Semi-supervised paradigm. Given a graph $G=(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges. $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix of the graph, $X\in\mathbb{R}^{N\times F}$ is the node feature matrix. The number of layers of the GNN model is $K$, the node representation in layer $k\in\{1,2,\ldots K\}$ is $z^{k}\in\mathbb{R}^{N\times H}$, where $H$ denotes the representation dimension in hidden layer. In this task, each node belongs to a specific category, only the labels of nodes in the training set are visible, and the target is to predict the category of nodes unlabeled.
Graph Attention Network. GAT is a graph neural network architecture that can adaptively learn the importance of neighboring nodes by leveraging attention mechanism to get the weights of neighbors [25]. The graph attention layer depicts how to get the representation $z^{k}\in\mathbb{R}^{N\times H_{k}}$ in layer $k$ with the input of representation $z^{k-1}\in\mathbb{R}^{N\times H_{k-1}}$ in layer $k-1$. The attention coefficient $\alpha_{ij}$ between a pair of nodes $(i,j)$ where $A_{ij}\neq 0$ is calculated by a linear transformation layer $W\in\mathbb{R}^{H_{k-1}\times H_{k}}$ and a shared attention mechanism $a:\mathbb{R}^{H_{k-1}}\times\mathbb{R}^{H_{k}}\rightarrow\mathbb{R}$ according to Eq.1.
@@
Where $\Delta(\cdot)$ stands for concatenation, average or other pooling operations.
Causal Inference. Causal inference is a new science for data [54], by making causal claims rather than merely associational claims with the belief that causality is inherently more stable. It is concerned with (1) Causal Discovery: Is there causal relationship between two variables? How does the cause impact the effect? And (2) Causal Effect Estimation: How much does the cause impact the effect? There are several important notations used in this paper.
  • Cause and Effect. A variable $X$ is said to be a cause of a variable $Y$ if $Y$ can change in response to changes in $X$. Or we can say $Y$ is ‘Listen to’ $X$. Then $X$ is the Cause and $Y$ is the Effect. If $Y$ directly responds to $X$, then $X$ is the Direct Cause of $Y$.
  • Causal Graph. A causal graph is a Directed Acycling Graph (DAG) that model the causality with graphical language. In Causal Graphs, every parent is a Direct Cause of all its children.
  • Intervention. If we do intervention in a variable $Z$ in the causal graph, it deletes all the edges from the parent variables and set the intervened variable to $\bar{z}$. We can denote it as the operation $do(Z=\bar{z})$. The children of $Z$ would change naturally with the change of $Z$.
  • Total Effect (TE). Total Effect measures the whole effect on $Y$ caused by $X$, including the direct effect and indirect effect. The TE can be counted by $TE_{x\rightarrow\bar{x}}=Y(X=x)-Y(do(X=\bar{x}))$.
With the Preliminaries we proposed above, we can depict the causal graph underlying GAT in accordance with Eq.1 and Eq.2, as shown in Figure 2.
@@
  • $V_{N}\rightarrow D\rightarrow\alpha_{i}$: When the normalization of attention coefficients is applied, the neighboring nodes will influence the attention distribution of the central node through the degree of the central node.
It is not difficult to notice that $V_{N}$ affects the final attention distribution $\alpha_{i}$ of $V_{i}$ through two causal paths. On the one hand, the representation of the $V_{N}$ itself affects its importance to $V_{i}$. On the other hand, the degree of $V_{i}$ changes due to the existence of $V_{j}\in V_{N}$, and thus affects the final attention coefficients distribution when normalizing $\alpha$. In order to measure the effect of one (or more) neighboring node(s) on the central node’s learned attention, we choose TE to carry out the calculation on causal effect of neighboring nodes, which can be adopted as a measurement of their DE.
We estimate the TE by intervening the LND of $V_{i}$. As illustrated in Figure 3, for a neighboring node $V_{j}\in V_{N}$, $V_{j}=0$ represents the reservation of $V_{j}$ as a neighbor of $V_{i}$, while $V_{j}=1$ represents removing $V_{j}$ from $V_{N}$. According to the Eq.3, we can display how to estimate the effect on the attention coefficient distribution of $V_{i}$ caused by $V_{j}$.
@@
Similarly, we denote the self-attention coefficient that $V_{j}$ assigns to itself as $\alpha_{self\_attention}$, and we can get the TE of $V_{j}$ for the self-attention of $V_{i}$ according to Eq.4:
@@
Figure 3: Intervention on the local neighbor distribution of the central node.
@@ 4 Pre-experiments and Key Hypothesis
In this section, we depict our observation from designed pre-experiments in Section 4.1 and key hypothesis in Section 4.2. In the pre-experiments, we disentangle the effect of neighboring nodes into two factors and manipulate them to generate intervened graphs as treatment groups. The experiment results indicate that nodes in the same class can provide similar semantic information for discrimination, where Class-level Semantic Space Hypothesis can be derived. We also propose the inference of Class-level Semantic Space Hypothesis, Low Distraction and High Self-attention, which is the core strategy of our method.
4.1 The Effect of Local Neighbor Distribution(LND)
The most important characteristic of GNN is the ability to aggregate the information provided by neighboring nodes and update the representation of the central node, therefore, the Local Neighbor Distribution (LND) is an important contributing factor on the ability of GNN models. As illustrated in Figure 4a, LND can be decomposed into two factors, Class-wise and Degree. The former statistically characterizes the distribution of neighboring nodes in different classes, and the latter denotes the number of neighboring nodes. Figure 4b illustrates that different Class-wise will change the local homophily of central node, and Figure 4c illustrates that LND with same homophily would be significantly different under different Degree. We formalize the LND of nodes in graph $G$ as $LND_{G}=\left\{W_{c},D_{c}\right\},c\in C$, where $C$ denotes the set of classes of nodes in $G$.
@@ 4.1 The Effect of Local Neighbor Distribution(LND)
To further discover the influence of LND on discrimination ability of GNN models, we set the graph $G$ as control group and intervene the LND of $G$ to get different treatment groups, for control experiments set up. Then we conduct control experiments on three representative GNN models, GCN [55], GraphSAGE [56] and GAT [25], and compare their node classification accuracy (the outcome of treatment, which is represented as $Y$) on different groups. The experiment setting is illustrated as Figure 5 and Table 1. We choose two heterophilic graph dataset, Chameleon and Squirrel to conduct the pre-experiment. The setting of experiments groups is as follwing:
@@
    Treatment group 1: reduce the Degree of the central node, and keep the Class-wise constant.
@@
    Treatment group 2: reduce the Degree of the central node, and set the Class-wise randomly. We set three random groups with the random seed 0, 10, 100.
@@
    Treatment group 3: the Degree of the central nodes remains constant, while randomly replace the neighboring nodes with different nodes of same class. We search the nodes to replace with the random seed 0, 10, 100.
Figure 5: The control and treatment group setting in the pre-experiment.
Table 1: Control experiment setting
@@
The result of control experiments are illustrated in Figure 6. It can be noted that:
@@
    $Y\left(t_{0}\right)\approx Y\left(t_{3}\right)$. It indicates that connections between different classes are substitutable, and keeping both $W$ and $D$ constant, but changing the nodes specifically connected in LND has little effect on the discrimination ability. It also indirectly indicates that nodes in the same class provide similar semantic information. We refer to this type of semantic information as Class-Level Semantic.
@@
    $Y\left(t_{1}\right)\approx Y\left(t_{2}\right)>Y\left(t_{0}\right)$. After removing a portion of the neighbors while keeping $W$ constant, it is instead able to improve the discrimination of graph nodes; while the improvement can be achieved by reducing $D$ and randomly alter $W$, which indicates that a number of connections in the graph are meaningless, and the distribution of such meaningless connecting edges does not present a difference in different classes.
@@
Figure 6: The result of the control experiments.
@@ 4.2 Low Distraction and High Self-attention
Drawing on the observations from pre-experiments, we propose a hypothesis and its inference. As illustrated in Figure 7, for a highly heterophilic graph $G$ with nodes of three different classes, the ideal underlying semantic space is comprised of three compact clusters, and each cluster is composed of the mapped graph nodes belong to this class. Each cluster has different size, density and location, and can be distinguished easily from each other. When we observe more nodes of each class, we can obtain a more accurate distribution and shape of its Semantic Clusters. Using a limit-thinking approach, if we could obtain all nodes of a certain class, the Semantic Cluster observed could represent the distribution of all nodes of that class in the semantic space. We refer to this as a Class-level Semantic Cluster, and the ideal space is referred to as Class-level Semantic Space.
Figure 7: An example portrait of ideal semantic space (dimension = 3). The circle with different color represents the distribution of Class-level Semantic Cluster of different class, and the quadrilateral star represents the center of each Class-level Semantic Cluster.
Hypothesis 1: Class-level Semantic Space Hypothesis. The graph $G$ is able to be mapped to an ideal $d$-dimension semantic space $S=f(G)$ , in which nodes of the same class situate very closely and nodes of different classes are as far away from each other as possible. Since an ideal Semantic Cluster is compact, the cluster center can serve as a representation of the distribution of nodes of that class in the semantic space. Therefore, the connections between different classes are substitutable, where each Semantic Cluster center is denoted as Eq.5:
@@
In the Class-level Semantic Space, it is evident that the closer the central node is to its Semantic Cluster center, the stronger its discrimination ability will be. Since the message passing mechanism aggregates features from neighboring nodes to the central node, neighbors from different classes exert a force that tends to push the central node away from its own Semantic Cluster center, which can be seen as a distraction and should be minimized. Conversely, both neighbors from the same class and the self-attention generate a force that pulls the central node closer to its own Semantic Cluster center, which should be reinforced. In a highly heterophilic graph where neighbors from the same class are rare, it is essential to enhance self-attention while mitigating distraction.
Inference 1: Low Distraction and High Self-attention. For nodes in a heterophilic graph, when nodes make more use of their own information and ignoring information from nodes of different classes while aggregating information, their final representation will be closer to the Semantic Cluster center of their classes in $S$.
Proof. For the central node $v_{i}$ and its neighboring node $v_{j}$, let the aggregation weight be $w$, we can get the representation of $v_{i}$ after aggregation and updating as $h_{i}=\sigma\left(w_{i}\cdot z_{i}+\sum\limits_{j\in N_{i}}w_{j}\cdot z_{j}\right)$, where $h_{i}$ is closer to $s_{ci}$, model’s discrimination ability on $v_{i}$ will be stronger. When the graph is highly heterophilic, $h_{i}$ can be represented as Eq.6:
@@
Because we hope $h_{i}$ is closer to $s_{ci}$, the optimization object is $\max\left(w_{i}+\sum\limits_{m}w_{m}-\sum\limits_{n}w_{n}\right)$. For the sake of condition in heterophilic graphs that $m\leq n$, we hope the weight for the central node itself $w_{i}$ can be maximized, equals to enhancing the self-attention and avoiding the distraction of dissimilar neighbors.
GAT, functioning as an implicit graph neural network architecture [57], is able to adaptively learn the weights of nodes for guiding the aggregation process. On the one hand, it may be easier to pose the distraction crisis to the central node due to the high proportion of inter-class edges on the heterophilic graph. On the other hand, by learning the weight distribution with Low distraction and High Self-attention, GAT can directly enhance its discrimination ability. Therefore, we foster strengths and circumvent weakness for GAT, leveraging the learned attention distribution as signal to guide the GAT identify and remove the Distraction Neighbors. The operation of graph trimming doesn’t require architecture alternation or new neighbors searching, but learns an optimal attention distribution by enhancing the self-attention.
@@ 5.1 The Architecture of CAT
The Causally graph Attention network for Trimming heterophilic graph (CAT) proposed in this paper mainly contains two important modules: Class-level Semantic Clustering Module and Total Effect Estimation Module. The former obtains the basic unit to estimate TE of neighboring nodes on central node’s attention learning, and the latter further estimates the TE by graph intervention. We introduced the CAT in Algorthim 1, where the $\Theta_{W}$ and $\Theta_{W2}$ of GAT represents the model parameters for learning feature transformation and attention distribution respectively. The pipeline of CAT is illustrated in Figure 8. As illustrated in Figure 8, the framework of CAT can adopt different GATs as the base model, and implement two modules to obtained the trimmed graph that can optimize the attention distribution of the base GAT.
@@
    Class-level Semantic Clustering Module. This module is derived from the Class-level Semantic Space Hypothesis, which maps the LND of the central node to a space that can discriminate class-level semantic better. The Semantic Clusters output in this module further serve as the basic object for estimating TE on the central node.
@@
    Total Effect Estimation Module. This module is derived from the Low distraction and High Self-attention, which obtains the TE of each class on the central node by the intervention of different Semantic Clusters. The Distraction Neighbors are identified in accordance with the TE and removed to get the final trimmed graph.

### v3

@@ 3 Preliminaries
Semi-supervised Graph Node Classification. Graph node classification is a fundamental task in graph representation learning, to classify graph nodes into predefined categories [39], and can be used as a proxy task for measuring the discrimination ability of graph representation models. Existing methods mainly focus on the semi-supervised paradigm. Given a graph $G=(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges. $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix of the graph, and $X\in\mathbb{R}^{N\times F}$ is the node feature matrix. The number of layers in the GNN model is $K$, and the node representation in layer $k\in\{1,2,\ldots K\}$ is $z^{k}\in\mathbb{R}^{N\times H}$, where $H$ denotes the representation dimension in the hidden layer. In this task, each node belongs to a specific category, only the labels of the nodes in the training set are visible, and the goal is to predict the category of unlabelled nodes.
Graph Attention Network. GAT is a graph neural network architecture that can adaptively learn the importance of neighboring nodes by leveraging an attention mechanism to obtain the weights of neighbors [3]. A graph attention layer depicts how to obtain the representation $z^{k}\in\mathbb{R}^{N\times H_{k}}$ in layer $k$ with the input of representation $z^{k-1}\in\mathbb{R}^{N\times H_{k-1}}$ in layer $k-1$. The attention coefficient $\alpha_{ij}$ between a pair of nodes $(i,j)$ where $A_{ij}\neq 0$ is calculated by a linear transformation layer $W\in\mathbb{R}^{H_{k-1}\times H_{k}}$ and a shared attention mechanism $a:\mathbb{R}^{H_{k-1}}\times\mathbb{R}^{H_{k}}\rightarrow\mathbb{R}$ according to Eq.1.
@@
Here $\Delta(\cdot)$ stands for concatenation, averaging or another pooling operation.
Causal Inference. Causal inference is a new data science [40], that involves making causal claims rather than merely associational claims based on the belief that causality is inherently more stable. It is concerned with (1) causal discovery (Is there a causal relationship between two variables? How does the cause impact the effect?) and (2) causal effect estimation (How much does the cause impact the effect?) Important notations used in this paper are as follows.
  • Cause and Effect. A variable $X$ is identified as a cause of a variable $Y$ if $Y$ can change in response to changes in $X$. Alternatively, we can say that $Y$ is ’Listen to’ $X$. Then $X$ is the cause and $Y$ is the effect. If $Y$ directly responds to $X$, then $X$ is the direct cause of $Y$.
  • Causal Graph. A causal graph is a Directed Acycling Graph (DAG) that models the causality with graphical language. In causal graphs, every parent is a direct cause of its children.
  • Intervention. If we intervene in a variable $Z$ in a causal graph, it deletes all the edges from its parent variables and sets the intervened variable to $\bar{z}$. We can denote this operation as $do(Z=\bar{z})$. The children of $Z$ change naturally with the change in $Z$.
  • Total Effect (TE). Total effect measures the whole effect of $X$ on $Y$, including the direct effect and indirect effect. The TE can be calculated by $TE_{x\rightarrow\bar{x}}=Y(X=x)-Y(do(X=\bar{x}))$.
With the preliminaries we stated above, we construct the causal graph underlying GATs in accordance with Eq.1 and Eq.2, as shown in Figure 2.
@@
  • $V_{N}\rightarrow D\rightarrow\alpha_{i}$: When the attention coefficients are normalized, the neighboring nodes influence the attention distribution of the central node through the degree of the central node.
Notably, $V_{N}$ affects the final attention distribution $\alpha_{i}$ of $V_{i}$ through two causal paths. On the one hand, the representation of $V_{N}$ affects its importance to $V_{i}$. On the other hand, the degree of $V_{i}$ changes due to the existence of $V_{j}\in V_{N}$, thereby influencing the final attention coefficient distribution when normalizing $\alpha$. To measure the effect of one (or more) neighboring node(s) on the learned attention of the central node, we choose TE to calculate the causal effect of neighboring nodes, which is adopted as measurements of their DE.
We estimate the TE by intervening in the LND of $V_{i}$. As illustrated in Figure 3, for a neighboring node $V_{j}\in V_{N}$, $V_{j}=0$ represents the reservation of $V_{j}$ as a neighbor of $V_{i}$, while $V_{j}=1$ represents the removal of $V_{j}$ from $V_{N}$. According to Eq.3, we can estimate the effect of $V_{j}$ on the attention coefficient distribution of $V_{i}$.
@@
Similarly, we denote the self-attention coefficient that $V_{j}$ assigns to itself as $\alpha_{self\_attention}$, and we can obtain the TE of $V_{j}$ on the self-attention of $V_{i}$ according to Eq.4:
@@
Figure 3: Intervention implemented for the local neighbor distribution of the central node.
@@ 4 Pre-experiments and Key Hypothesis
In this section, we illustrate our observations derived from the pre-experiments designed in Section 4.1 and propose the key hypothesis in Section 4.2. In the pre-experiments, we disentangle the effects of neighboring nodes into two factors and intervene in them to generate different intervention graphs as treatment groups. The experimental results indicate that nodes in the same class can provide similar semantic information for discrimination, where Class-level Semantic Space Hypothesis can be derived. We also propose the inference of Class-level Semantic Space Hypothesis, Low Distraction and High Self-attention, which is the core strategy of our method.
4.1 The Effect of the Local Neighbor Distribution(LND)
GNNs are renowned for the ability to aggregate the information provided by neighboring nodes and update the representation of the central node. Therefore, the local neighbor distribution (LND) is an important contributing factor to the ability of GNN models. As illustrated in Figure 4a, the LND can be decomposed into two factors, Class-wise ($W$) and Degree ($D$). The former statistically characterizes the distribution of neighboring nodes in different classes, and the latter denotes the number of neighboring nodes. Figure 4b illustrates that different $W$ will change the local homophily of the central node, and Figure 4c illustrates that LNDs with the same homophily are significantly different under different $D$. We formalize the LND of nodes in graph $G$ as $LND_{G}=\left\{W_{c},D_{c}\right\},c\in C$, where $C$ denotes the set of node classes in $G$.
@@ 4.1 The Effect of the Local Neighbor Distribution(LND)
To further determine the influence of the LND on the discrimination ability of GNNs, we used $G$ as a control group and intervened in the LND of $G$ to construct different treatment groups. Then, we conduct control experiments on three representative GNN models, GCN [41], GraphSAGE [42] and GAT [3], and compare their node classification accuracies (the outcomes, which are represented as $Y$) on different groups. The experimental settings are illustrated in Figure 5 and Table 1. We choose two heterophilic graph datasets, Chameleon and Squirrel to conduct the pre-experiment. The experimental settings are as follows:
@@
    Treatment Group 1: The $D$ of the central node decreases, and the $W$ remains constant.
@@
    Treatment Group 2: The $D$ of the central node decreases, and the $W$ is set randomly. We set three random groups with random seeds of 0, 10, and 100.
@@
    Treatment Group 3: The $D$ of the central nodes remains constant, while the neighboring nodes are randomly replaced with different nodes belonging to the same class. We search for replacement nodes with random seeds 0, 10, and 100.
Figure 5: The control and treatment group settings used in the pre-experiment.
Table 1: Settings for control experiments.
@@
The results of the control experiments are illustrated in Figure 6. The following can be noted:
@@
    $Y\left(t_{0}\right)\approx Y\left(t_{3}\right)$. This indicates that the connections between different classes are substitutable, and both $W$ and $D$ are held constant. However, changing the nodes specifically connected in the LND has little effect on the discrimination ability of the GNN. It also indirectly indicates that nodes in the same class provide similar semantic information. We refer to this type of semantic information as Class-Level Semantic.
@@
    $Y\left(t_{1}\right)\approx Y\left(t_{2}\right)>Y\left(t_{0}\right)$. After removing a portion of the neighbors while keeping $W$ constant, the GNN can better discriminate graph nodes; This improvement can be achieved by reducing $D$ and randomly altering $W$, which indicates that some connections in the graph are meaningless, and the distribution of such meaningless connecting edges does not vary with different classes.
@@
Figure 6: The results of the control experiments.
@@ 4.2 Low Distraction and High Self-attention
Drawing on the observations from the pre-experiments, we propose a hypothesis and its inference. As illustrated in Figure 7, for a highly heterophilic graph $G$ with nodes belonging to three classes, the ideal semantic space is composed of three compact clusters, and each cluster is composed of the mapped graph nodes belonging to their associated class. Each cluster has unique size, density, and location parameters, and the clusters can be easily distinguished from others. Observing more nodes of each class contributes to a more accurate distribution of its Semantic Clusters. Using a limit-thinking approach, if all nodes of a certain class are available, the Semantic Cluster observed can represent the distribution of all nodes belonging to that class. We refer to this as a Class-level Semantic Cluster, and the ideal space is referred to as Class-level Semantic Space.
Figure 7: An example of an ideal semantic space (dimension = 3). Circles with different colors represent the distribution of the Class-level Semantic Cluster of different classes, and the quadrilateral star represents the center of each Class-level Semantic Cluster. The arrows in the representation space indicate the forces acting between nodes, whose strength is determined by the attention allocated by the central node.
Hypothesis 1: Class-level Semantic Space Hypothesis. A graph $G$ can be mapped to an ideal $d$-dimensional semantic space $S=f(G)$, where belonging to the same class are located very close and nodes of different classes are as far away as possible. Since an ideal Semantic Cluster is compact, the cluster center can serve as the representation of nodes belonging to that class in the semantic space. Therefore, the connections between different classes are substitutable, where each Semantic Cluster center is denoted as Eq.5:
@@
In the Class-level Semantic Space, it is evident that the closer the central node is to its Semantic Cluster center, the stronger its discrimination ability will be. Since the message passing mechanism aggregates features from neighboring nodes to the central node, neighbors from different classes exert a force that pushes the central node away from its own Semantic Cluster center, which is a distraction and should be minimized. Conversely, both neighbors from the same class and self-attention generate forces that pull the central node closer to its own Semantic Cluster center, which should be reinforced. In a highly heterophilic graph with few same-class neighbors, it is essential to enhance self-attention to mitigate distraction.
Inference 1: Low Distraction and High Self-attention. When a node in heterophilic graphs makes more use of its own information and ignores information derived from nodes of different classes during aggregation, its final representation will be closer to the Semantic Cluster center of its class in $S$.
Proof. For the central node $v_{i}$ and its neighboring node $v_{j}$, letting the aggregation weight be $w$, we can obtain the representation of $v_{i}$ after the aggregation and updating process as $h_{i}=\sigma\left(w_{i}\cdot z_{i}+\sum\limits_{j\in N_{i}}w_{j}\cdot z_{j}\right)$, where $h_{i}$ is closer to $s_{ci}$, model’s discrimination ability for $v_{i}$ will be stronger. When the graph is highly heterophilic, $h_{i}$ can be represented as Eq.6:
@@
Because we hope $h_{i}$ is closer to $s_{ci}$, the optimization target is $\max\left(w_{i}+\sum\limits_{m}w_{m}-\sum\limits_{n}w_{n}\right)$. For the sake of the heterophilic graph condition stating that $m\leq n$, we hope that the weight of the central node itself $w_{i}$ can be maximized, which is equal to enhancing the self-attention level and avoiding the distraction caused by dissimilar neighbors.
GAT adaptively learns the weights of nodes to guide the aggregation. On the one hand, it may be easier to pose the distraction crisis to the central node due to the high proportion of interclass edges in a heterophilic graph. On the other hand, by learning a weight distribution with Low distraction and High Self-attention, GAT can directly enhance its discrimination ability. Therefore, we foster strengths and circumvent weaknesses for GAT by leveraging the learned attention distribution as signals, to guide the GAT to identify and remove the Distraction Neighbors. The graph trimming operation does not require architecture alternations or new neighbor searches but rather learns an optimal attention distribution to enhance self-attention.
@@ 5.1 The Architecture of CAT
The Causal graph Attention network for Trimming heterophilic graphs (CAT) proposed in this paper mainly contains two important modules: the Class-level Semantic Clustering Module and the Total Effect Estimation Module. The former obtains the basic unit for estimating the TE of the neighboring nodes, and the latter further estimates the TE via graph intervention. We introduced the CAT in Algorthim 1, where the $\Theta_{W}$ and $\Theta_{W2}$ of the GAT represent the model parameters for feature transformation and attention distribution learning, respectively. The pipeline of CAT is illustrated in Figure 8. As illustrated in Figure 8, the framework of CAT can adopt different GATs as the base model, and finally obtain a trimmed graph that can optimize the attention distribution of the base GAT.
@@
    Class-level Semantic Clustering Module. This module is derived from the Class-level Semantic Space Hypothesis, which maps the LND of the central node to a space that can better discriminate class-level semantics. The Semantic Clusters output in this module further serve as the basic object for estimating TE.
@@
    Total Effect Estimation Module. This module is derived from the Low distraction and High Self-attention, which obtains the TE of each class on the central node by intervening in different Semantic Clusters. The Distraction Neighbors are identified in accordance with the TE and removed to obtain the final trimmed graph.